OpenStudy (anonymous):

Assume that y varies inversely with x. If y = -3 when x = 2.4, what is y when x = 2? A)-3.6 B)-0.8 C)-7.2 D)4.8

OpenStudy (marziajackson$):

I would say c

OpenStudy (johnweldon1993):

\[\large y = \frac{k}{x}\] y = -3 when x = 2.4 so \[\large -3 = \frac{k}{2.4} \rightarrow k = -7.2\] If k = -7.2 we then have \[\large y = \frac{-7.2}{x}\] so now when x = 2...what is y? well we lug that 2 in for x and solve for 'y' \[\large y = \frac{-7.2}{2}\]